# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

### v2

@@ Appendix D Computational Complexity in R
Appendix E Limitations in Experiments
@@
During our experiment, we found that on large dense graphs, especially the one with small diameters such as ogbn-arxiv ($|V|=169k$, $diam=25$), it is difficult to compute our Total Influence measurement, as $\bar{N}_{H}$ in Equation (13) quickly converges to $N$ as $H$ increases. In this case, if the model also happens to have complex architectures (e.g., GraphGPS), the computation of $R$ and $T_{h}$ will therefore become impractical. We believe this constitutes a limitation of our proposed measurement, where future works may explore more effective methods for dense graphs based on stochastic sampling.
Appendix F Proofs and Justifications in Section 5
F.1 Details of Linearization
@@
F.2 Proofs for Theoretical Results in Section 5
Lemma F.1 (Eigenvalue complementarity of normalized operators).
@@ F.2 Proofs for Theoretical Results in Section 5
Proposition F.2 (Self-loops decrease algebraic connectivity of the original graph, from Section 5.2).
@@ F.3 Rationale in Theorem 5
Appendix G Jacobian-based Influence Score Motivation
Here, we provide additional motivation behind the proposed measurement for quantifying long-rangeness in Section 4.
G.1 Intuition behind Jacobian-based Influence Score
The Jacobian has been used in analysis of node interactions in GNNs in multiple previous works (Xu et al., 2018; Gasteiger et al., 2022; Di Giovanni et al., 2023). For example, it is used in Di Giovanni et al. (2023, Theorem 4.1) to show when over-squashing happens in long-range interactions, and to show how vanishing gradients occur in very deep GNNs. Influence specifically has been used to compute a natural measure of interactions between two nodes (Xu et al., 2018). We accordingly use aggregated influence, Equation (2), to gauge how nodes at a distance $h$ affect the output of the GNN at a focal node, thus quantifying long-rangeness. By the definition of partial derivatives, we can understand the Jacobian as follows:
@@ G.1 Intuition behind Jacobian-based Influence Score
G.2 Potential Limitations: Output Cancellation
@@
Proposition G.1 (Absolute Jacobian sensitivity may over-estimate influence).
@@
Definition G.2 (Message-Passing Neural Network layer).
@@
Definition G.3 (Smooth Hypersurface in $\mathbb{R}^{D}$).
@@
Theorem G.4 (Measure-zero of exact cancellation at MPNN initialization).
@@
defines a level set of smooth functions, a hypersurface (or a set of lower-dimensional submanifolds) in parameter space. The solution set of a nontrivial smooth function has Lebesgue measure zero unless it is identically zero across all $\bm{\theta}$, which is not the case here. Furthermore, since $\bm{\theta}$ is drawn from an absolutely continuous distribution (such as Gaussian or uniform), the probability of exactly selecting a parameter that lies on this hypersurface is zero at initialization. Thus, exact cancellation of Jacobian contributions is an event of measure zero in the space of our idealized single-layer MPNN: this can naturally be extended to multiple layers. Note that in this proof we have assumed the Jacobian sum is not identically zero by construction, unlike in Proposition G.1 where the function was explicitly constructed to ensure cancellation. ∎
Lastly, it is worth noting an alternative perspective on interpreting influence measures derived from the sum of absolute Jacobian entries, particularly when considering potential cancellation effects as demonstrated in Proposition G.1. One can indeed argue that the measure’s utility lies precisely in its capacity to quantify the magnitude of sensitivity to inputs from individual distant nodes or pathways, irrespective of whether these influences ultimately negate one another in contributing to the final output. From this stance, focusing on the sum of absolute values reveals the underlying structure and strength of potential long-range dependencies (the information channels themselves) that might be obscured by observing only the net effect. This interpretation, therefore, hinges on defining dependency or interaction based on the existence and intensity of these information flow pathways, rather than strictly on their final, combined impact on a node’s prediction.
Appendix H Additional Definitions in Section 6
Definition H.1 (Standard lattice in $\mathbb{R}^{D}$).
@@ Appendix H Additional Definitions in Section 6
Definition H.3 (Quasi-isometric graphs).
@@
Appendix I Proofs for Theoretical Results in Section 6

### v3

@@ Appendix E Computational Complexity
Appendix F Limitations and Future Work
Total Influence under model bias.
Since our measurement is based on the gradient of the underlying model, its behavior will be naturally influenced by the bias in the architecture. For example, as described in Appendix D.3, when the models’ depth (num_layers=64) is much deeper than the known task depth ($k=16$), we observe that $R$s are generally larger than those from models with 16 layers in our city networks. Meanwhile, the per-hop influence also suggests that deeper models leverage information from distant hops beyond the ground truth range $k$ (which makes sense since the model has no information about $k$, and node features beyond the $k$th hop may contain useful information for prediction). Therefore, the biased approximation of the ground-truth function will inevitably lead to a biased estimation of the underlying task’s range, even though the measurement remains a faithful description of how the model utilizes long-range information.
However, our conclusions regarding long-rangeness across datasets are still valid, since each model is validated across different datasets under the same depth in our analysis, where all models show clear long–range patterns on our city-networks compared to those on the existing benchmarks at depth=16 (Section 4) and depth=64 (Appendix D.3).
@@ Appendix F Limitations and Future Work
During our experiment, we found that on large dense graphs, especially the one with small diameters such as ogbn-arxiv ($|V|=169k$, $diam=25$), it is difficult to compute our Total Influence measurement, as $\bar{N}_{H}$ in equation 13 quickly converges to $N$ as $H$ increases. In this case, if the model also happens to have complex architectures (e.g., GraphGPS), the computation of $R$ and $T_{h}$ will therefore become impractical. Meanwhile, we also want to point out that this is a common challenge for all Jacobian-based analysis in the literature (Xu et al., 2018; Gasteiger et al., 2022), and it is an open question for future works to explore more effective methods for measuring long-range dependency on dense graphs.
Inductive setting.
In the current work, we focus on testing long-range signals under transductive settings on large-scale graphs, which is largely underexplored in the literature. However, we believe it is also possible to extend our method to inductive settings by sampling a set of cities’ road networks via OpenStreetMap (or similar geographic graph sources), and then defining a graph-level classification or regression task on them, e.g., urban (graph) morphology classification. We leave this direction for future explorations.
Appendix G Proofs and Justifications in Section 5
G.1 Details of Linearization
@@
G.2 Proofs for Theoretical Results in Section 5
Lemma G.1 (Eigenvalue complementarity of normalized operators).
@@ G.2 Proofs for Theoretical Results in Section 5
Proposition G.2 (Self-loops decrease algebraic connectivity of the original graph, from Section 5.2).
@@ G.3 Rationale behind Theorem 5
G.4 Relations to Braess Paradox and Previous Work
Sparsity of the graph and spectral gap.
Our main theoretical result in Section 5 (Theorem 5.2) suggests that the second largest eigenvalue of the normalized augmented adjacency matrix has a lower bound that is increasing in graph diameter and decreasing in maximum node degree, hence the spectral gap tends to be smaller for graphs with larger diameter and smaller maximum degree. However, this tendency may not be exact as the bound is not necessarily tight, and furthermore, a smaller maximum degree does not necessarily imply a sparser graph (although the two can be generally correlated). Therefore, our result is not in contradiction with the Braess paradox, which is an interesting observation by itself.
Spectral gap and tendency of over-smoothing.
Our analysis of the rate of over-smoothing is purely based on the spectral analysis that the speed of convergence to a stationary point upon repeated application of a matrix operator depends on the spectral gap: a smaller gap generally slows down convergence in this precise sense. It is worth noting that this analysis is task-agnostic, i.e., it does not take into account node classification as a task, and furthermore distribution of node labels. The analysis of over-smoothing by Jamadandi et al. (2024) is, however, task-dependent, as it specifically highlights situations when pruning edges can mitigate over-smoothing of task-beneficial signals by disconnecting nodes with different labels. Therefore, while both are meaningful, the two analyses look at over-smoothing from slightly different perspectives.
Appendix H Motivation for Jacobian-based Influence Score
We provide additional motivation behind the proposed measurement for quantifying long-rangeness in Section 4.
H.1 Intuition behind Jacobian-based Influence Score
The Jacobian has been used in analysis of node interactions in GNNs in multiple previous works (Xu et al., 2018; Gasteiger et al., 2022; Di Giovanni et al., 2023). For example, it is used in Di Giovanni et al. (2023, Theorem 4.1) to show when over-squashing happens in long-range interactions, and to show how vanishing gradients occur in very deep GNNs. Influence specifically has been used to compute a natural measure of interactions between two nodes (Xu et al., 2018). We accordingly use aggregated influence, Equation equation 2, to gauge how nodes at a distance $h$ affect the output of the GNN at a focal node, thus quantifying long-rangeness. By the definition of partial derivatives, we can understand the Jacobian as follows:
@@ H.1 Intuition behind Jacobian-based Influence Score
H.2 Potential Limitations: Output Cancellation
@@
Proposition H.1 (Absolute Jacobian sensitivity may over-estimate influence).
@@
Definition H.2 (Message-Passing Neural Network layer).
@@
Definition H.3 (Smooth Hypersurface in $\mathbb{R}^{D}$).
@@
Theorem H.4 (Measure-zero of exact cancellation at MPNN initialization).
@@
defines a level set of smooth functions, a hypersurface (or a set of lower-dimensional submanifolds) in parameter space. The solution set of a nontrivial smooth function has Lebesgue measure zero unless it is identically zero across all $\bm{\theta}$, which is not the case here. Furthermore, since $\bm{\theta}$ is drawn from an absolutely continuous distribution (such as Gaussian or uniform), the probability of exactly selecting a parameter that lies on this hypersurface is zero at initialization. Thus, exact cancellation of Jacobian contributions is an event of measure zero in the space of our idealized single-layer MPNN: this can naturally be extended to multiple layers. Note that in this proof we have assumed the Jacobian sum is not identically zero by construction, unlike in Proposition H.1 where the function was explicitly constructed to ensure cancellation. ∎
Lastly, it is worth noting an alternative perspective on interpreting influence measures derived from the sum of absolute Jacobian entries, particularly when considering potential cancellation effects as demonstrated in Proposition H.1. One can indeed argue that the measure’s utility lies precisely in its capacity to quantify the magnitude of sensitivity to inputs from individual distant nodes or pathways, irrespective of whether these influences ultimately negate one another in contributing to the final output. From this stance, focusing on the sum of absolute values reveals the underlying structure and strength of potential long-range dependencies (the information channels themselves) that might be obscured by observing only the net effect. This interpretation, therefore, hinges on defining dependency or interaction based on the existence and intensity of these information flow pathways, rather than strictly on their final, combined impact on a node’s prediction.
Appendix I Additional Definitions in Section 6
Definition I.1 (Standard lattice in $\mathbb{R}^{D}$).
@@ Appendix I Additional Definitions in Section 6
Definition I.3 (Quasi-isometric graphs).
@@
Appendix J Proofs for Theoretical Results in Section 6